OpenStudy (anonymous):

What is the value of the expression 1 over 3 to the negative 5 power

OpenStudy (anonymous):

@keke_luvs_u

OpenStudy (anonymous):

@Darry&Gator

OpenStudy (anonymous):

Do you mean...\[\frac{ 1 }{ 3 }^{-5}\]

OpenStudy (anonymous):

\[\left( \frac{ 1 }{ 3 } \right)^{-5}=\frac{ 1 }{ \left( \frac{ 1 }{ 3 } \right)^{5} }=\frac{ 1 }{ \left( \frac{ 1 }{3^5 } \right)}=1\times \frac{ 3^5 } { 1 }\] \[=3^5\]

OpenStudy (anonymous):

@Darry&Gator

OpenStudy (anonymous):

\[\frac{ 1 }{ 3^{-5} }=3^5\]

OpenStudy (anonymous):

Well that explains it. (^_^) Good job @surjithayer !